# JUHAN ARU, PHILÉMON BORDEREAU

### v1

@@ 1.3. Dirichlet energies
In what follows, we will often be brought to consider diverging Dirichlet energies of harmonic functions that are locally constant on the boundary of a domain. Consider $D$ a regular bounded simply connected domain again with $x\neq y\in\partial D$, with boundary $\mathcal{B}_{l}\cup\mathcal{B}_{r}=(yx)\cup(xy)$ as in the notation section. Let $u_{D}$ denote the harmonic extension of the boundary values $-1$ on $\mathcal{B}_{l}$, $+1$ on $\mathcal{B}_{r}$. The Dirichlet energy $\norm{u_{D}}_{\nabla(D)}:=\norm{\nabla u_{D}}_{D}$ is clearly ill-defined, but we can regularize it at the discontinuity points of $u_{D}$:
$\norm{u_{D}}_{\nabla(D),\text{reg}}^{2}:=\lim_{\varepsilon_{1}\to 0,\varepsilon_{2}\to 0}\int_{z\in D,\absolutevalue{z-x}>\varepsilon_{1},\absolutevalue{z-y}>\varepsilon_{2}}\absolutevalue{\nabla u_{D}}^{2}\mathrm{d}z+\frac{4}{\pi}(\log\varepsilon_{1}+\log\varepsilon_{2}).$
@@
It is easy to see that the limit exists. Nonetheless, in many relevant situations below, we will consider differences of such diverging Dirichlet energies to which it is possible to provide a meaning without renormalization. Consider in the above set-up $D^{\prime}\subset D$ another test domain such that $\mathrm{dist}(\mathcal{B}_{l},D\setminus D^{\prime})>0$. One might formally write
$\int_{D}\absolutevalue{\nabla u_{D}}^{2}$ $=-\int_{\mathcal{B}_{l}}\partial_{\nu}u_{D},$
$\int_{D^{\prime}}\absolutevalue{\nabla u_{D^{\prime}}}^{2}$ $=-\int_{\mathcal{B}_{l}}\partial_{\nu}u_{D^{\prime}},$
and define
$\norm{u_{D^{\prime}}}_{\nabla(D^{\prime})}-\norm{u_{D}}_{\nabla(D)}:=\int_{\mathcal{B}_{l}}\partial_{\nu}(u_{D}-u_{D^{\prime}})$
@@
where the normal derivative is well-defined and the integral is finite since $u_{D}-u_{D^{\prime}}$ is harmonic in $D^{\prime}$ with boundary values 00 on $\partial D\cap\partial D^{\prime}$ and $u_{D}-1\geq 0$ on $\delta$. One can also replace $u_{D},u_{D^{\prime}}$ by their expression involving the Poisson kernel:
$u_{D}(\cdot)=1-\int_{\mathcal{B}_{l}}H_{D}(\cdot,\mathrm{d}y),\quad u_{D^{\prime}}(\cdot)=1-\int_{\mathcal{B}_{l}}H_{D^{\prime}}(\cdot,\mathrm{d}y)$
so that the definition above becomes (using that $\partial_{\nu_{z}}H_{D}(z,w)=-H_{\partial D}(z,w)$ here, with $\nu_{z}$ the outwards normal vector)
$\norm{u_{D^{\prime}}}_{\nabla(D^{\prime})}-\norm{u_{D}}_{\nabla(D)}$ $:=\iint_{\mathcal{B}_{l}\times\mathcal{B}_{l}}(H_{\partial D}-H_{\partial D^{\prime}})(\mathrm{d}z,\mathrm{d}w)$
$=(1_{\mathcal{B}_{l}},(H_{\partial D}-H_{\partial D^{\prime}})1_{\mathcal{B}_{l}})_{\mathcal{B}_{l}}.$
@@
This also corresponds to the mass of Brownian excursions from $\mathcal{B}_{l}$ that hit $D\setminus D^{\prime}$. In particular, the resulting expression is conformally invariant. The difference of these Dirichlet energies can also be seen as the difference of the inverse extremal distance $M(\mathcal{B}_{l},\mathcal{B}_{r}^{\prime})-M(\mathcal{B}_{l},\mathcal{B}_{r})$, see [Ahl10, Theorem 4-5]. The formula above also clearly generalizes when $D$ is a general admissible domain (possibly multiply connected), and $\partial D=\mathcal{B}_{-}\cup\mathcal{B}_{+}$ with $\mathcal{B}_{-}$, $\mathcal{B}_{+}$ both containing an arc from $x$ to $y$, as well as the other boundary components of $\partial D$.
1.4. $\mathrm{SLE}$ measures in simply connected domains
@@
Let us recall the construction of $\mathrm{SLE}_{\kappa}$ in the upper half-plane, before reviewing them as non-probability measures. Given $\kappa>0$, let $B_{\kappa\cdot}$ be a one-dimensional Brownian motion running at speed $\kappa$, and $(g_{t}(z))_{t\geq 0}$ be the solution to the following chordal Loewner equation:
@@
For each $z\in\mathbb{H}$, the flow above is well-defined until $T(z)=\inf\left\{t\geq 0:g_{t}(z)-B_{\kappa t}=0\right\}$. The resulting hulls $K_{t}:=\left\{z\in\mathbb{H}:T(z)\leq t\right\}$ almost surely correspond to simple curves for $0\leq\kappa\leq 4$, and are generated by non-crossing curves for $4<\kappa$ (i.e. there exists $\gamma$ such that $K_{t}$ is equal to the unbounded connected component of $\mathbb{H}\setminus\gamma_{t}$), as proven in [RS05]. The probability measure on these curves generated by the Brownian motion is called the $\mathrm{SLE}_{\kappa}$ measure in $\mathbb{H}$ from 00. Furthermore, $\gamma(t)\to\infty$ almost surely for every $\kappa>0$, which allows to define $\mathrm{SLE}_{\kappa}$ from $x$ to $y$ in $D$ a domain with $x\neq y\in\partial D$ as the image of $\mathrm{SLE}_{\kappa}$ in $\mathbb{H}$ (from 00 to $\infty$) by a conformal map $\varphi:\mathbb{H}\to D,\varphi(0)=x,\varphi(+\infty)=y$. We denote these measures $\nu_{D,\kappa}(z,w)$, probability measures for the moment. Because there remains freedom in precomposing $\varphi$ by a dilation, we will without further mention see $\nu_{D,\kappa}(z,w)$ as measures on curves up to (positive) time reparametrization.
@@
The one-parameter family of measures $\left\{\nu_{D,\kappa}(z,w)\right\}_{(D,z,w)}$ in fact describes the only probability measures on curves from $z$ to $w$ in $D$ satisfying conformal invariance (clear from the construction) and a conformal Markov property, that states that the conditional law of $\gamma([t,+\infty))$ under $\nu_{D,\kappa}(z,w)$ given $\gamma_{t}$ is equal to $\nu_{D\setminus\gamma_{t},\kappa}(\gamma(t),w)$ (see [Sch00]). From now on we make the dependence on $\kappa$ implicit in the notation, and restrict to the range $\kappa\leq 4$ where the curves are simple.
@@
In multiply connected domains, the two properties above are not sufficient to characterize $\mathrm{SLE}$ [^1] For instance, one could define a different SLE type of curve in a conformal annulus, that would behave like $\mathrm{SLE}_{2}$ when the conformal modulus is greater than one, like $\mathrm{SLE}_{4}$ otherwise. The conformal invariance requirement is too weak capture this change of behaviour of the curve because it cannot compare the law of the curve for different conformal classes of domains.. However, these curves satisfy an extra property, called the restriction property or boundary invariance rule [LSW03, Law09]: for $K\subset\mathbb{H}$ a compact hull such that $\mathrm{dist}(K,0)>0$ and $\mathbb{H}\setminus K$ is connected, it holds that $\nu_{\mathbb{H}\setminus K}(0,+\infty)\ll\nu_{\mathbb{H}}(0,+\infty)$, and that
(5) $\frac{\mathrm{d}\nu_{\mathbb{H}\setminus K}(0,+\infty)}{\mathrm{d}\nu_{\mathbb{H}}(0,+\infty)}(\eta)=\mathbf{1}_{\eta\subset\mathbb{H}\setminus K}\varphi^{\prime}(0)^{-h}\exp(\frac{\mathbf{c}}{2}m_{\mathbb{H}}(\eta,K)),$
@@
where $\varphi:\mathbb{H}\setminus K\to\mathbb{H}$ is a conformal equivalence such that $\varphi(0)=0$ and $\varphi(z)=z+o(1)$ as $z\to\infty$. Notice that the term $\varphi^{\prime}(0)$ is positive and corresponds to the probability that a Brownian excursion from 00 to $\infty$ in $\mathbb{H}$ does not hit $K$ [LSW03]: similarly, for $D$ a simply connected (non-crossing) admissible domain and $D^{\prime}$ a test domain,
@@
where $H_{\partial D^{\prime}}(x,y)/H_{\partial D}(x,y)$ is again the probability that an excursion from $x$ to $y$ in $D$ remains in $D^{\prime}$. It was suggested in [Law09] to use this extra property to characterize $\mathrm{SLE}$. Moreover, we now no longer see $\mathrm{SLE}$ as probability measures, but rather as measures with mass $Z_{D}(x,y)=Z_{D,\kappa}(x,y)$. Note that some regularity of $\partial D$ at $x$ and $y$ is required to define this mass or partition function: but this is not restrictive at the level of the corresponding probability measures, as any (finitely connected) domain is conformally equivalent to one with analytic boundary around the images of $x$ and $y$, so that the $\mathrm{SLE}$ probability measure can be defined there and brought back to the initial domain.
Definition 3.
$\mathrm{SLE}$ is the unique family of measures $\left\{\nu_{D}(x,y)\right\}_{(D,x,y)}$ on curves (modulo time-reparametrization) in $D$ a general admissible domain from $x$ to $y$ of total mass $Z_{D}(x,y)=\norm{\nu_{D}(x,y)}$ normalized so that $Z_{\mathbb{H}}(0,+\infty)=1$, satisfying the following:
@@
    (7) $\varphi_{\ast}\nu_{D}(x,y)=\absolutevalue{\varphi^{\prime}(x)}^{h}\absolutevalue{\varphi^{\prime}(y)}^{h}\nu_{D^{\prime}}(f(x),f(y)).$
@@
  • Boundary perturbation Let $D^{\prime}\subset D$ be admissible domains agreeing in a neighborhood of $x$ and $y$. Then $\nu_{D^{\prime}}(x,y)\ll\nu_{D}(x,y)$, with Radon-Nikodym derivative given by
@@
Note that in the case of simply connected domains, it is consistent with Equations (5) and (6) to define $Z_{D}(x,y)=H_{D}(x,y)^{b}$ when $x,y\neq\infty$, and $Z_{D}(x,y)=\absolutevalue{\varphi^{\prime}(x)}^{b}$ when $y=\infty$, where $\varphi:D\to\mathbb{H}$ is a conformal equivalence with $\varphi(x)=0$ and $\varphi(z)=z+o(1)$ as $z\to\infty$.
In [Law11], these measures are constructed in any general admissible domain, from a Carathéodory argument: the restriction of $\nu_{D}(x,y)$ to curves contained in a general test domain $D^{\prime}$ is prescribed by the boundary perturbation rule, and the restriction to different test domains are compatible thanks to the additive structure of the loop measure. This justifies the denomination of ‘test domains’: to identify a measure on curves in a domain $D$ as the $\mathrm{SLE}$ measure it is sufficient to verify Equation (8) for $D^{\prime}\subset D$ simply connected. Note that uniqueness also follows from the construction.
We further stress that some conditions in the definition above are redundant: the domain Markov property for $\mathrm{SLE}$ in general domains is a consequence of the fact that this property holds on simply connected domains, and is inherited through the boundary perturbation rule (provided that the partition function is finite: see again [Law11] for all the details). Reversibility of the $\mathrm{SLE}$ is also inherited from the simply connected case, which was proven in [Zha08].
On the other hand, finiteness of the partition function is less clear. In [Law11], it is shown to be finite when $\kappa\leq 8/3$, or when $D$ is a conformal annulus. In this paper, we construct explicitly the $\mathrm{SLE}_{4}$ measure and the $\mathrm{SLE}_{\kappa}$ measure for $8/3<\kappa\leq 4$ in the non-crossing case, and provide an expression for their partition function that proves finiteness.
2. $\mathrm{SLE}_{4}$ as a level-line of Gaussian free fields
@@ 2.1.1. Gaussian free field
$\mathbb{E}\left[\Phi_{f}\Phi_{g}\right]=\iint_{D\times D}f(x)G_{D}(x,y)g(y)\mathrm{d}x\mathrm{d}y<+\infty,$
where $G_{D}$ is the Green function, previously introduced. It is possible to define the GFF as a random element of $H^{0-}(D)$ tested against $H_{0}^{1}(D)$ (see for instance [WP21, Dub09]), so that instead of $\Phi_{f}$ we can write $(\Phi,f)$ to denote the distributional pairing. For our purposes it will suffice to see the GFF as an element of $H^{-1}(D)$. For $g$ a harmonic function in $D$, the GFF with boundary condition $g$ is understood as $\Phi^{g}:=\Phi^{0}+g$, where $\Phi^{0}$ is a zero-boundary GFF in $D$, and from now on we let $\Phi_{D}^{g}$ be our notation for a GFF in $D$ with mean $g$.
For $g\in H_{0}^{1}(D)$, $\Phi_{D}^{g}$ is also a zero-boundary GFF under the law that is also absolutely continuous to that of $\Phi_{D}^{0}$, with Radon-Nikodym derivative given by
$\exp((\Phi_{D}^{0},g)_{\nabla(D)}-(g,g)_{\nabla(D)}),$
where $(\cdot,\cdot)_{\nabla(D)}$ is the Dirichlet inner product and $(\Phi_{D}^{0},g)_{\nabla(D)}:=(\Phi_{D}^{0},-\Delta g)$.  [^2] Note that it is possible to naturally make sense of the GFF tested against functions in $H^{-1}(D)$, such as $-\Delta g\in H^{-1}(D)$: see [BP24]. This in fact only holds when $g\in H_{0}^{1}(D)$, identifying the latter as the Cameron-Martin space of the GFF.
The GFF also enjoys a natural domain Markov property. For $D$ a domain and $\delta$ a smooth crosscut splitting $D$ in $D_{1}$ and $D_{2}$, we can write
$\Phi_{D}^{0}=\Phi_{D_{1}}^{0}+u_{\Phi_{D}^{0}|_{\delta}}+\Phi_{D_{2}}^{0}$
where $u_{\Phi_{D}^{0}|_{\delta}}$ is a random harmonic function in $D\setminus\delta$, and all summands in the right-hand side are independent. In fact, one can define (see [Dub09]) a pathwise trace of $\Phi_{D}^{0}$ on $\delta$ as $\Phi_{D}^{0}|_{\delta}\in H^{-1}(\delta)$, whose law is a centered Gaussian process of covariance matrix $G_{D}|_{\delta\times\delta}$, so that $u_{\Phi_{D}^{0}|_{\delta}}$ is the harmonic extension of $\Phi_{D}^{0}|_{\delta}$ to $D\setminus\delta$ (which is indeed harmonic as the Poisson kernel is smoothing).
Note that the domain Markov property for the GFF holds in a much more general setting [WP21], although one may no longer interpret the random harmonic function $h$ as the extension of the values of the GFF on some compact set.
We are now interested in the study of random subsets of $D$, coupled with the GFF, for which a version of the Markov property holds.
Definition 4 (Local sets).
@@
Changing the measure with a shift of the GFF given by some $g\in H_{0}^{1}(D)$ as in the Cameron-Martin theorem is also compatible with local sets, as shows the following lemma from [ALS22]:
Lemma 5 (Lemma 3.1 in [ALS22]).
Consider $(\Phi:=\Phi_{D}^{0},A,\Phi_{A})$ a local set coupling, $\tilde{\Phi}=\Phi-g$, and $g_{A}$ the orthogonal projection of $g$ in $H_{0}^{1}(D)$ on the subspace of functions that are harmonic in $D\setminus A$. Set $\tilde{\Phi}_{A}:=\Phi_{A}-g_{A}$. Then $(\tilde{\Phi},A,\tilde{\Phi}_{A})$ is a local set coupling and the law of $(A,\Phi_{A})$ under $\Phi$ is absolutely continuous with respect to the law of $(A,\tilde{\Phi}_{A}+g_{A})=(A,\Phi_{A})$ under $\tilde{\Phi}$, with Radon-Nikodym derivative given by
$\exp((\Phi_{A},g)_{\nabla}-\frac{1}{2}(g_{A},g_{A})_{\nabla}).$
Lastly, we will later need to compute the expectation of expressions similar to that appearing in the previous lemma, for $g$ a Gaussian random variable. The following lemma provides a formula that will be relevant in this computation.
Lemma 6 (Proposition 1.2.8 in [DPZ02]).
Let $H$ be a Hilbert space with $Q$ a positive, trace class operator, $M$ symmetric such that $Q^{1/2}MQ^{1/2}<I$, $m\in H$. Denote by $N_{Q}$ the centered Gaussian measure on $H$ with covariance $Q$. Then

### v2

@@ 1.3. Dirichlet energies
We will often consider diverging Dirichlet energies of harmonic functions that are locally constant on the boundary of a domain. Consider $D$ a smooth bounded simply connected domain again with distinct $x,y\in\partial D$ and boundary $\mathcal{B}_{l}\cup\mathcal{B}_{r}=(yx)\cup(xy)$ as in the notation section. Let $u_{D}$ denote the harmonic extension of the boundary values 1-1 on $\mathcal{B}_{l}$, +1+1 on $\mathcal{B}_{r}$. The Dirichlet energy $\norm{u_{D}}_{\nabla(D)}:=\norm{\nabla u_{D}}_{D}$ diverges, but we can regularise it at the discontinuity points of $u_{D}$:
[…]
It is easy to see that the limit exists. Furthermore, it can be verified (see for instance [12], proof of Proposition 5.2) that the regularised Dirichlet energy satisfies the following conformal covariance property, for $D$ a smooth bounded simply connected domain $\varphi:D\to\varphi(D)$ a bounded conformal equivalence:
[…]
so that up to a multiplicative constant,
(8) $\exp(\norm{u_{D}}_{\nabla(D),\mathrm{reg}}^{2})=H_{\partial D}(x,y)^{-4/\pi}.$
@@
Nonetheless, in many relevant situations below, we will consider differences of such diverging Dirichlet energies to which it is possible to provide a meaning without renormalisation. Consider in the above set-up $D^{\prime}\subset D$ another test domain such that $\mathrm{dist}(\mathcal{B}_{l},D\setminus D^{\prime})>0$. One might formally write
[…]
with $\partial_{\nu}$ the inwards normal derivative, and define
$\norm{u_{D^{\prime}}}_{\nabla(D^{\prime})}^{2}-\norm{u_{D}}_{\nabla(D)}^{2}:=2\int_{\mathcal{B}_{l}}\partial_{\nu}(u_{D^{\prime}}-u_{D})$
@@
where the normal derivative is well-defined and the integral is finite since $u_{D^{\prime}}-u_{D}$ is harmonic in $D^{\prime}$ with boundary values 0 on $\mathcal{B}_{l}$ and $1-u_{D}\geq 0$ on $\partial D^{\prime}\setminus\mathcal{B}_{l}$. This is of course equal to the difference of the regularised Dirichlet energies. One can also replace $u_{D},u_{D^{\prime}}$ by their expression involving the Poisson kernel:
[…]
so that the definition above becomes
[…]
$=4(1_{\mathcal{B}_{l}},(H_{\partial D}-H_{\partial D^{\prime}})1_{\mathcal{B}_{l}})_{\mathcal{B}_{l}}.$
@@
This also corresponds to (four times) the mass of Brownian excursions from $\mathcal{B}_{l}$ that hit $D\setminus D^{\prime}$. In particular, the resulting expression is conformally invariant. The difference of these Dirichlet energies can also be seen as (four times) the difference of the inverse extremal distance $\mathrm{ED}(\mathcal{B}_{l},\mathcal{B}_{r}^{\prime})^{-1}-\mathrm{ED}(\mathcal{B}_{l},\mathcal{B}_{r})^{-1}$, see [1, Theorem 4-5]. The formula above applies also to a general admissible domain $D$ (possibly multiply connected), where $\partial D=\mathcal{B}_{-}\cup\mathcal{B}_{+}$ with $\mathcal{B}_{-}$, $\mathcal{B}_{+}$ containing resp. $(yx)$ and $(xy)$, as well as the other boundary components of $\partial D$.
1.4. Conformal restriction SLE
We start by recalling the construction of chordal $\mathrm{SLE}_{\kappa}$ in the upper half-plane and then present Lawler’s definition of what we call conformal restriction SLE.
1.4.1. SLE in simply-connected domains
@@
Given $\kappa>0$, let $B_{\kappa\cdot}$ be a one-dimensional Brownian motion running at speed $\kappa$, and $(g_{t}(z))_{t\geq 0}$ be the solution to the following chordal Loewner equation:
@@
For each $z\in\mathbb{H}$, the flow above is well-defined until $T(z)=\inf\left\{t\geq 0:g_{t}(z)-B_{\kappa t}=0\right\}$.
@@
The resulting hulls $K_{t}:=\left\{z\in\mathbb{H}:T(z)\leq t\right\}$ almost surely correspond to simple curves for $0\leq\kappa\leq 4$, and are generated by non-crossing curves for $4<\kappa$ (i.e. there exists $\gamma$ such that $K_{t}$ is equal to the unbounded connected component of $\mathbb{H}\setminus\gamma_{t}$), as proved in [33, Theorem 5.1, 6.1]. Furthermore, $\gamma(t)\to\infty$ almost surely for every $\kappa>0$. The probability measure on these curves generated by the Brownian motion is called the chordal $\mathrm{SLE}_{\kappa}$ measure in $\mathbb{H}$ from 0 to $\infty$ and it is scale invariant.
$\mathrm{SLE}_{\kappa}$ from $x$ to $y$ in a simply-connected domain $D$ with distinct $x,y\in\partial D$ is then defined as the image of $\mathrm{SLE}_{\kappa}$ in $\mathbb{H}$ (from 0 to $\infty$) by a conformal map $\varphi:\mathbb{H}\to D,\varphi(0)=x,\varphi(+\infty)=y$. We denote these probability measures $\nu_{D,\kappa}(z,w)$ and from now on understand them as measures on curves up to (positive) time reparametrisation (since the conformal map $\varphi$ above is not unique, but there remains freedom in precomposing by a dilation).
@@
The one-parameter family of measures $\left\{\nu_{D,\kappa}(z,w)\right\}_{(D,z,w)}$ in fact describes the only probability measures on curves from $z$ to $w$ in $D$ satisfying conformal invariance (clear from the construction) and a conformal Markov property, that states that the conditional law of $\gamma([t,+\infty))$ under $\nu_{D,\kappa}(z,w)$ given $\gamma_{t}$ is equal to $\nu_{D\setminus\gamma_{t},\kappa}(\gamma(t),w)$ (see [36]). For the rest, for readability let us leave out the dependence on $\kappa$ in the notation, and restrict ourselves to the range $\kappa\leq 4$ where the curves are simple.
1.4.2. Conformal restriction SLE
@@
In multiply connected domains, the two properties above are not sufficient to characterise $\mathrm{SLE}$ [^2] For instance, one could define a different SLE type of curve in a conformal annulus, that would behave like $\mathrm{SLE}_{2}$ when the conformal modulus is greater than one and like $\mathrm{SLE}_{4}$ otherwise. The conformal invariance requirement is too weak capture this change of behaviour of the curve because it cannot compare the law of the curve for different conformal classes of domains.. However, in the simply connected case, these curves satisfy an extra property, called the restriction property or boundary invariance rule [26] (see also [24, Proposition 2.1]): for $K\subset\mathbb{H}$ a compact hull such that $\mathrm{dist}(K,0)>0$ and $\mathbb{H}\setminus K$ is (connected and) simply connected, it holds that $\nu_{\mathbb{H}\setminus K}(0,+\infty)\ll\nu_{\mathbb{H}}(0,+\infty)$ and that
[…]
@@
where $\varphi:\mathbb{H}\setminus K\to\mathbb{H}$ is the conformal equivalence such that $\varphi(0)=0$ and $\varphi(z)=z+o(1)$ as $z\to\infty$. We note that the term $|\absolutevalue{\varphi^{\prime}(0)}$ corresponds to the probability that a Brownian excursion from 0 to $\infty$ in $\mathbb{H}$ does not hit $K$ [26, Proposition 4.1]: similarly, for $D$ a simply connected admissible domain and $D^{\prime}$ a test domain,
@@
where $H_{\partial D^{\prime}}(x,y)/H_{\partial D}(x,y)$ is again the probability that an excursion from $x$ to $y$ in $D$ remains in $D^{\prime}$. It was suggested in [24] to use this property to characterise $\mathrm{SLE}$ in general domains, provided that we consider $\mathrm{SLE}$ as measures with mass $Z_{D}(x,y)=Z_{D,\kappa}(x,y)$. Note that some regularity of $\partial D$ at $x$ and $y$ is required to define this mass or partition function: but this is not restrictive at the level of the corresponding probability measures, as any (finitely connected) domain is conformally equivalent to one with smooth boundary around the images of $x$ and $y$, so that the $\mathrm{SLE}$ probability measure can be defined there and brought back to the initial domain.
Definition 1.3 (Conformal restriction SLE, [25]).
$\mathrm{SLE}$ is the unique family of measures $\left\{\nu_{D}(x,y)\right\}_{(D,x,y)}$ on curves (modulo time-reparametrisation) in $D$ a general admissible domain from $x$ to $y$ of total mass $Z_{D}(x,y)=\norm{\nu_{D}(x,y)}$ normalised so that $Z_{\mathbb{H}}(0,+\infty)=1$, satisfying the following:
@@
[…]
@@
  • Boundary perturbation/restriction property Let $D^{\prime}\subset D$ be admissible domains agreeing in a neighbourhood of $x$ and $y$. Then $\nu_{D^{\prime}}(x,y)\ll\nu_{D}(x,y)$, with Radon-Nikodym derivative given by
@@
[…]
In [25, Section 4], these measures are constructed in any general admissible domain, from a Carathéodory argument: the restriction of $\nu_{D}(x,y)$ to curves contained in a general test domain $D^{\prime}$ is prescribed by the boundary perturbation rule, and the restriction to different test domains are compatible thanks to the additive structure of the loop measure. This justifies the denomination of ‘test domains’: to identify a measure on curves in a domain $D$ as the $\mathrm{SLE}$ measure it is sufficient to verify Equation (12) for $D^{\prime}\subset D$ simply connected. As a matter of fact, it is enough to take $D^{\prime}$ with smooth boundary, as the identity then holds for all subdomains by approximation (see [26] for more details about approximations of hulls by smooth hulls). Note that uniqueness also follows from the construction.
We further stress that some conditions in the definition above are redundant: the domain Markov property and conformal covariance for $\mathrm{SLE}$ in general domains is a consequence of the fact that these properties hold on simply connected domains, and are inherited through the boundary perturbation rule (provided that the partition function is finite: see again [25, Section 4] for all the details). Reversibility of the $\mathrm{SLE}$ is also inherited from the simply connected case, which was proved in [47].
On the other hand, finiteness of the total mass i.e. the partition function is less clear. In [25], it is shown to be finite when $\kappa\leq 8/3$, or when $D$ is a conformal annulus. In this paper, we construct explicitly the $\mathrm{SLE}_{4}$ measure in all topological cases and the $\mathrm{SLE}_{\kappa}$ measure for $8/3<\kappa\leq 4$ in the non-crossing case, and provide an expression for their partition function that proves finiteness.
2. $\mathrm{SLE}_{4}$ as a mixture of level-lines of Gaussian free fields
@@ 2.1.1. Gaussian free field
$\mathbb{E}\left[\Phi_{f_{1}}\Phi_{f_{2}}\right]=\iint_{D\times D}f_{1}(x)G_{D}(x,y)f_{2}(y)\mathrm{d}A(x)\mathrm{d}A(y)<+\infty,$
where $G_{D}$ is the Green function previously introduced. It is possible to define the GFF as a random element of $H^{-\varepsilon}(D)$ for any $\varepsilon>0$ [43, 12]); in particular, this justifies the writing of $\Phi_{f}$ as $(\Phi,f)_{D}$. For our purposes it will suffice to see the GFF as an element of $H^{-1}(D)$. For $g$ a harmonic function in $D$, the GFF with boundary condition $g$ is understood as $\Phi^{g}:=\Phi^{0}+g$, where $\Phi^{0}$ is a zero-boundary GFF in $D$. From now on we let $\Phi_{D}^{g}$ be our notation for a GFF in $D$ with mean $g$.
For $f\in H_{0}^{1}(D)$, $\Phi_{D}^{0}+f$ is also a zero-boundary GFF under the law that is also absolutely continuous w.r.t. that of $\Phi_{D}^{0}$, with Radon-Nikodym derivative given by
$\exp((\Phi_{D}^{0},f)_{\nabla(D)}-(f,f)_{\nabla(D)}),$
where $(\cdot,\cdot)_{\nabla(D)}$ is the Dirichlet inner product and $(\Phi_{D}^{0},f)_{\nabla(D)}:=(\Phi_{D}^{0},-\Delta f)$.  [^3] Note that it is possible to naturally make sense of the GFF tested against functions in $H^{-1}(D)$, such as $\Delta f\in H^{-1}(D)$: see [8, Section 1.8]. This in fact only holds when $f\in H_{0}^{1}(D)$, identifying the latter as the Cameron-Martin space of the GFF.
The GFF also enjoys a natural domain Markov property [43, Chapter 4]. We will use it here in the following set-up. For $D$ a domain and $\delta$ a smooth crosscut splitting $D$ in $D_{1}$ and $D_{2}$, we can write
$\Phi_{D}^{0}=\Phi_{D_{1}}^{0}+v_{\Phi_{D}^{0}|_{\delta}}+\Phi_{D_{2}}^{0}$
where $v_{\Phi_{D}^{0}|_{\delta}}$ is a random harmonic function in $D\setminus\delta$, and all summands on the right-hand side are independent. In fact, one can define (see [12, Section 4.3.2]) a pathwise trace of $\Phi_{D}^{0}$ on $\delta$ as $\Phi_{D}^{0}|_{\delta}\in H^{-1}(\delta)$, whose law is a centred Gaussian process of covariance $G_{D}|_{\delta\times\delta}$, so that $v_{\Phi_{D}^{0}|_{\delta}}$ defines the harmonic extension of $\Phi_{D}^{0}|_{\delta}$ to $D\setminus\delta$ (which is indeed harmonic as the Poisson kernel is smoothing).
An important class of random subsets of $D$, commonly called local sets, are those with which the GFF satisfies a strong Markov property.
Definition 2.1 (Local sets).
@@
Changing the measure with a shift of the GFF given by some $f\in H_{0}^{1}(D)$ as in the Cameron-Martin theorem is also compatible with local sets, as shows the following lemma from [3]:
Lemma 2.2 (Lemma 3.1 in [3]).
Consider $(\Phi:=\Phi_{D}^{0},A,\Phi_{A})$ a local set coupling, $\tilde{\Phi}=\Phi-f$, and $f_{A}$ the orthogonal projection of $f$ in $H_{0}^{1}(D)$ on the subspace of functions that are harmonic in $D\setminus A$. Set $\tilde{\Phi}_{A}:=\Phi_{A}-f_{A}$. Then $(\tilde{\Phi},A,\tilde{\Phi}_{A})$ is a local set coupling and the law of $(A,\Phi_{A})$ under $\Phi$ is absolutely continuous with respect to the law of $(A,\tilde{\Phi}_{A}+f_{A})=(A,\Phi_{A})$ under $\tilde{\Phi}$, with Radon-Nikodym derivative given by
$\exp((\Phi_{A},f)_{\nabla}-\frac{1}{2}(f_{A},f_{A})_{\nabla}).$
Lastly, we will later need to compute the expectation of expressions similar to that appearing in the previous lemma, for $f$ a random function, normally distributed. The following lemma provides a formula that will be relevant in this computation.
Lemma 2.3 (Proposition 1.2.8 in [11]).
Let $H$ be a Hilbert space with $Q$ a positive, trace class operator, $M$ symmetric such that $Q^{1/2}MQ^{1/2}<I$, $m\in H$. Denote by $N_{Q}$ the centred Gaussian measure on $H$ with covariance $Q$. Then